OpenStudy (anonymous):

What is the oblique asymptote of the funtion -13x3-22x2-86x-50/x2+x+6?

OpenStudy (anonymous):

\[\LARGE {-13x^3-22x^2-86x-50 \over x^2+x+6}\] this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok.. I'll ask you a simple question: do you know how to find asymptotes? how about this? can you find obliques asymptote of this? \[\LARGE f(x)={x^2\over x+2}\] ?

OpenStudy (anonymous):

The only thing i know is you divide the domain so my first guess would be that i know the answer begins with x but the +2 confuses me.

OpenStudy (anonymous):

Ok.. If I solve MY example, step by step, would you try to solve yours, following its steps ? :)

OpenStudy (anonymous):

Yes, that would be helpful

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

so we have: \[\LARGE f(x)={x^2 \over x+2 }\] now we should use this equation: \[\LARGE y=kx+b\quad \quad ,\quad k\neq 0\] and we have: \[\Large k=\lim_{x\to \infty}{f(x)\over x}\quad \quad \text{and}\quad \quad b=\lim_{x\to \infty} \left[{f(x)-kx}\right]\] so \[\Large k=\lim_{x\to \infty}{f(x)\over x}=\lim_{x\to \infty} \frac{\frac{x^2}{x+2}}{x}= \] \[\Large =\lim_{x\to \infty} \frac{x^2}{x(x+2)}=\lim_{x\to \infty}\frac{x}{x\left(1+\frac2x\right)} =\] \[\Large =\frac{1}{1+0}=1\] so k=1 now we find b \[\Large b=\lim_{x\to \infty}[f(x)-kx]=\lim_{x\to\infty}[\frac{x^2}{x+2}-1\cdot x]=\] \[\Large =\lim_{x\to \infty}\left[\frac{\cancel{x^2}\cancel{-x^2}-2x}{x+2}\right]=...=-2\] so b=-2 we substitute y=kx+b y=1*x-2 y=x-2

OpenStudy (anonymous):

Oh wow. Thank you very much for your help. I think once I keep reviewing the help you have given me I will be able to figure the problem out. I'm taking an online class and it's difficult without a teacher actually standing in front of you, but I think I'll understand it eventually lol. Thank you again.

OpenStudy (anonymous):

My pleasure, I used a "simplier" function just to avoid big numbers, ;) Good Luck.

OpenStudy (anonymous):

and remember one thing. \[\LARGE k\neq 0\] so.. when you'll be doing it, if you'll get k=0 don't continue any further. That's it. There's no oblique asymptote :)... Take care my friend ...

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

:)